Calculate and draw to scale. Mary crawls 5m north, then crawls 8m west, then crawls 12m east. Draw the resultant displacement of Mary.
Answer #1

20+ north 15 10+ В. IA о west east theta O -20 - 15 - 10 -5 UT 10 15 20 -5 10 15 south 20

Mary goes from O to A and then A to B and then B to C.

so resultant dsiplacement is OC

now OC = sq. root of (4^2+5^2) = 6.40 m

now tan theta = 5/4 so theta = 51.34 degress

so resultant dsiplacement is 6.40 m, 51.3 degrees north of east